### v1

# for Machine-Learned Interatomic Potentials

Abstract
Machine-learned interatomic potentials (MLIPs) are deployed for high-throughput materials screening without formal reliability guarantees. We show that a single MLIP used as a stability filter misses 93% of density functional theory (DFT)-stable materials (recall 0.07) on a 25,000-material benchmark. Proof-Carrying Materials (PCM) closes this gap through three stages: adversarial falsification across compositional space, bootstrap envelope refinement with 95% confidence intervals, and Lean 4 formal certification. Auditing CHGNet, TensorNet and MACE reveals architecture-specific blind spots with near-zero pairwise error correlations (; ), confirmed by independent Quantum ESPRESSO validation (20/20 converged; median DFT/CHGNet force ratio ). A risk model trained on PCM-discovered features predicts failures on unseen materials (area under the ROC curve [AUC-ROC] ) and transfers across architectures (cross-MLIP AUC-ROC ; feature importance ). In a thermoelectric screening case study, PCM-audited protocols discover 62 additional stable materials missed by single-MLIP screening—a 25% improvement in discovery yield.
Keywords: machine-learned interatomic potentials, adversarial testing, formal verification, Lean 4, materials discovery, uncertainty quantification
1 Introduction
Universal MLIPs—CHGNet1, MACE4, TensorNet27, ALIGNN35, SevenNet36, EquiformerV237—underpin high-throughput materials screening38, yet they are deployed without formal reliability guarantees. Matbench Discovery2 and earlier benchmarks10 evaluate 45+ models on 257K WBM (Wang–Botti–Marques3) structures, but aggregate accuracy metrics cannot answer the deployment-critical question: on which chemistries is this MLIP unreliable?
This creates a quantifiable safety gap. A single-MLIP stability screen achieves precision 0.47 and recall 0.07 across 25,000 WBM materials, missing 93.0% of DFT-stable candidates. The gap is not merely statistical: CHGNet rejects TlBiSe2 (a topological insulator with over 1,000 citations26) and Cs2KTlBr6 (a lead-free perovskite solar cell candidate with 1.27 eV band gap)—precisely the materials that high-throughput pipelines aim to find. Aggregate benchmarks showing “near-DFT accuracy” obscure these consequential, chemistry-specific blind spots.
We reframe MLIP reliability as a falsifiable safety claim in the sense of Dalrymple et al.’s Guaranteed Safe AI13: a world model (the MLIP), a safety specification (bounds on acceptable error), and a verifier (machine-checked proofs)45. This mirrors proof-carrying code15, where untrusted programs ship with proofs of safety.
Proof-Carrying Materials (PCM) instantiates this in three stages (Fig. 1):
1.
Adversarial falsification—automated adversaries (six strategies: random, heuristic, grid, LHS, Sobol, LLM) probe for failure regions in compositional space;
2.
Envelope refinement—counterexamples tighten the safety claim into bounds with bootstrap 95% confidence intervals (CIs);
3.
Formal certification—the refined envelope compiles into Lean 4 proofs with explicit physical axioms.
The framework is oracle-agnostic: swapping the MLIP requires no code changes.
Our central findings are threefold. First, MLIP blind spots are architecture-specific: three MLIPs produce near-zero pairwise force correlations (, , ; ) and fail on largely disjoint chemistries. Second, perturbation-based uncertainty quantification does not predict these compositional failures (, ), meaning structural UQ and adversarial compositional auditing capture independent failure dimensions. Third, and most importantly, PCM enables prospective validation: compositional features discovered through adversarial auditing predict failures on unseen materials with AUC-ROC and P@20% = 1.000 (perfect precision among the top-risk quintile), transforming retrospective audit into predictive intervention. These failure patterns transfer across architectures: a CHGNet-trained risk model predicts MACE failures (AUC-ROC = average cross-transfer) with feature importance correlation . Together, these results establish that (a) no single benchmark score captures reliability, (b) multi-MLIP auditing is essential for deployment, (c) formal verification combined with prospective prediction constitutes a new paradigm for MLIP validation, and (d) the full audit costs under $20.
2 Related Work
Adversarial MLIP testing. CAGO11 generates adversarial structures via uncertainty-calibrated geometry optimization on individual materials. PCM operates in compositional space—searching for chemical families where MLIPs systematically diverge from DFT. These are complementary: CAGO identifies which atomic arrangements are unreliable; PCM identifies which chemistries are unreliable regardless of arrangement. In our evaluation, perturbation-based uncertainty is not predictive of compositional failure (, ).
Uncertainty quantification. Conformal prediction19 provides distribution-free marginal coverage guarantees applied to interatomic potentials20. Monte Carlo dropout29, deep ensembles30 and chemical-space uncertainty estimation31; 32 have been applied to molecular property prediction. Critically, these methods provide marginal coverage (“95% of all materials fall within the prediction interval”) but cannot answer the deployment question which specific compositions are unreliable and why. Conformal prediction intervals widen uniformly across compositional space; they do not identify the compositional drivers (e.g. heavy elements, large unit cells) that concentrate failures, nor do they enable targeted DFT allocation for high-risk chemistries. PCM concentrates adversarial sampling on failure boundaries, producing tighter envelopes where they matter most: in a split conformal baseline on the same 5,000 WBM materials (2,500 calibration, 2,500 test), the feature-space conformal envelope at contains 1,924 materials (precision 6.8%) versus PCM’s adversary-informed bootstrap envelope with 97 materials (precision 7.2%)—a reduction in envelope size at comparable precision. Active learning12; 50; 51 selects informative points for retraining; PCM selects adversarial points for auditing—the adversarial budget (200 queries) is orders of magnitude smaller than retraining data requirements.
Safety cases and formal verification. AMLAS16 provides structured safety-case methodology for ML components but relies on human argumentation. Seshia et al.17 identify the specification gap as a central challenge for verified AI. Neural network verification43; 44 focuses on input–output properties of classifiers; PCM verifies domain-specific reliability claims about regression models. Adversarial testing in ML33; 34 targets input perturbations; falsification of cyber-physical systems7; 8 provides compositional testing frameworks. PCM instantiates the Guaranteed Safe AI framework13 for materials: the MLIP is the world model, the envelope is the specification, and Lean 4 proofs are the verifier output—providing machine-checkable evidence that structured safety cases currently lack.
3 Results
3.1 Architecture-specific failure profiles
We evaluate three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1)27; 28, and MACE-MP medium4—on 5,000 WBM materials. Structures are synthetic compositional probes—atoms placed at irrational (golden-ratio) fractional coordinates to avoid symmetry artifacts, with DFT-derived cell volumes (Methods)—designed to test whether each MLIP has learned the correct energy landscape for a given chemistry. The blind-spot classification itself—which materials are DFT-stable but MLIP-unstable—derives from the WBM benchmark, where both DFT and MLIPs evaluate fully relaxed equilibrium structures. Force failure threshold: 50 eV Å-1 (Fig. 2).
CHGNet fails on 31.1% of compositions (1,553/5,000, force 50 eV Å-1).
TensorNet fails on 75.7% (3,786/5,000).
MACE fails on 73.2% (3,659/5,000), producing catastrophically high forces ( 1,000 eV Å-1) on multiple compositions where CHGNet predicts moderate values.
All three pairwise force correlations are near zero: CHGNet–TensorNet [0.07, 0.13]; CHGNet–MACE [0.05, 0.21]; TensorNet–MACE [-0.02, -0.003] (bootstrap 95% CIs, ).
Failures are largely disjoint: 218 materials fail CHGNet but not MACE, while 2,324 fail MACE but not CHGNet; 337 fail TensorNet but not MACE, while 210 fail MACE but not TensorNet.
This is not a marginal difference. Three MLIPs, all trained on Materials Project data, produce qualitatively different force predictions on the same compositions. The higher absolute failure rates for TensorNet (75.7%) and MACE (73.2%) relative to CHGNet (31.1%) likely reflect differences in how each architecture generalises to far-from-equilibrium probe geometries; the scientifically important finding is not the absolute rates but the near-zero pairwise correlations and disjoint failure profiles, which demonstrate that architecture and training pipeline—not training data alone—determine which chemistries appear problematic. A stability screen using any single MLIP inherits that model’s specific blind spots—including several materials with known functional importance (Section 3.3).
3.2 Adversarial strategy comparison
Because the WBM–CHGNet benchmark has a 93.2% base counterexample (CX) rate—randomly sampling any composition has a chance of finding a failure—raw CX rate cannot distinguish adversary quality. We therefore evaluate six strategy types and seven LLM configurations (budget = 200) primarily on discovery diversity: the number of unique, previously unseen failure compositions found (Fig. 3, Extended Data Table 1).
Among automated strategies, all achieve high CX rates (86–89% for heuristic, random, and Sobol). LLM adversaries—tested as one strategy class among several—achieve marginally higher rates (88–100%), with their primary advantage being qualitative: convergence on high-, multi-element regions of compositional space (, ), discovering 36 DFT-stable/CHGNet-unstable materials not found by any baseline strategy. A shuffled-feature ablation confirms this works even with anonymous feature labels (88% CX vs 90% with real names; Extended Data Fig. 7), showing that the adversary–oracle separation—not materials expertise—is the core mechanism.
The full multi-adversary audit costs $18.13; the most cost-effective single strategy (Gemini 3 Flash as LLM adversary) discovers 17 unique materials for $0.05, though purely algorithmic strategies (Sobol, random) discover more unique materials overall (138 and 123, respectively) at zero API cost.
3.3 Functionally important discoveries
From the combined adversary outputs (Section 3.2), we rank materials by anomaly score (Eq. 1) and identify 21 unique materials with disjoint failure profiles (Table 1). Five contain actinides (Pu, Np, Th, U, Dy), three contain thallium—both underrepresented in CHGNet’s MPtrj training data.
TlBiSe2 is an experimentally verified topological insulator26 that CHGNet predicts as unstable while DFT confirms stability. Independent DFT recomputation (Quantum ESPRESSO, PAW-PBE) converges cleanly in 13 SCF steps while CHGNet produces 11.1 eV atom-1 energy and 19.9 eV Å-1 forces on the same golden-ratio probe structure—a catastrophic disagreement on a material with over 1,000 citations.
Cs2KTlBr6 is a lead-free halide double perovskite with a near-ideal 1.27 eV band gap for photovoltaics—precisely the class of non-toxic solar cell materials that high-throughput screening pipelines aim to find.
Five of 12 contain actinides relevant to nuclear fuel cycle modelling. If CHGNet were used as a stability filter, it would reject materials currently under active investigation in nuclear safety, topological electronics and photovoltaics.
| Rank | Formula | Anomaly | Pred. error (eV atom-1) | Max force (eV Å-1) | Application |
| 1 | DyLi2PuF8 | 3.49 | 7.06 | 75.6 | Nuclear fuel (molten salt) |
| 2 | Pb4Pt4O12 | 2.97 | 4.13 | 107.6 | Pt-group catalysis |
| 3 | Au2Ba2Np2Se6 | 2.70 | 6.95 | 44.9 | Nuclear waste |
| 4 | Au2Cs2Sc2Te6 | 2.68 | 5.36 | 73.4 | Thermoelectric |
| 5 | Bi2Cu2O4Te2V | 2.58 | 5.69 | 63.0 | Multiferroic |
| 6 | Th2I6 | 2.53 | 6.07 | 53.8 | Nuclear fuel processing |
| 7 | Rb2SnTlCl6 | 2.36 | 5.94 | 49.3 | Lead-free perovskite |
| 8 | Cs2KTlBr6 | 2.15 | 6.12 | 37.2 | Perovskite solar cell |
| 9 | Ba6NPb2Se | 1.61 | 4.96 | 36.0 | Optoelectronic |
| 10 | CsAuTe | 1.57 | 4.38 | 45.1 | Relativistic bonding |
| 11 | TlBiSe2 | 1.17 | 3.93 | 36.7 | Topological insulator |
| 12 | Cs2USe3 | 0.69 | 3.00 | 33.7 | Actinide chalcogenide |
3.4 Independent DFT validation at scale
To confirm that adversarially discovered blind spots are genuine—not artifacts of approximate structure construction—we recompute DFT ground states for the top 20 materials using Quantum ESPRESSO (pw.x v6.7, PAW-PBE, SSSP efficiency pseudopotentials, 60/480 Ry cutoffs, -mesh). Materials are selected from the WBM benchmark by: DFT-stable, CHGNet-unstable, max , , sorted by prediction error descending. Structures use golden-ratio fractional coordinate spacing with DFT-derived cell volumes (Methods).
All 20 materials converge to well-defined self-consistent field (SCF) ground states (11–150 iterations, 22–6,360 s per material on 8 CPU cores). Two materials (Ba2Pd4Sb4 and BaF6In) require expanded cells and tighter mixing parameters on retry, but ultimately converge. The converged materials span 27 elements across (Li) to (Ba). We compare forces rather than energies because DFT total energies include core electrons (e.g. eV/atom for Cu7Zn1) and are not directly comparable to CHGNet’s formation-energy-referenced predictions. Forces on identical structures provide a direct, reference-free comparison.
Across the 18 materials evaluated on identical golden-ratio structures, DFT forces exceed CHGNet forces by a median factor of (range: –; Extended Data Table Extended Data Table 2; force comparison excludes 2 retry materials whose structures differ). The headline result: Cu7Zn1 (brass, the most widely used copper alloy) converges in 15 SCF steps with DFT forces of 557 eV Å-1—while CHGNet predicts 36 eV Å-1, a underestimate. Cu7Mn1 (manganese bronze) shows a underestimate. These are not exotic compositions: they are industrial alloys that CHGNet’s training data should cover.
We emphasise that the blind-spot classification (DFT-stable, CHGNet-unstable) originates from the WBM benchmark, where both DFT and CHGNet evaluate relaxed equilibrium structures—not the synthetic probes used here. The golden-ratio structures serve only as a compositional probe: they test whether DFT and CHGNet agree on force magnitudes for a given chemistry. A model that has learned correct interatomic interactions for a composition should produce forces of comparable magnitude on any structure of that composition, not only near-equilibrium ones. The median force underestimate therefore indicates that CHGNet has not learned the correct energy landscape for these chemistries, consistent with their absence or underrepresentation in training data.
The 100% convergence rate (20/20), achieved on adversarially selected worst-case compositions, confirms that the PCM pipeline discovers genuine MLIP failures, not structure-generation artifacts.
As a complementary test, we relax all 20 DFT-validated adversarial discoveries using CHGNet v0.3.0 with BFGS optimization (fmax = 0.05 eV Å-1, max 500 steps). If the force disagreements were merely a consequence of probing far-from-equilibrium geometries, CHGNet relaxation should converge rapidly to low-energy minima. Instead, all 20 materials exhibit massive relaxation energy changes of 3.98–15.45 eV/atom—far exceeding the 0.1 eV/atom scale of typical formation energies. Eighteen materials converge (10–477 ionic steps), but two—Ag6Cl2S2 and As2In4Pd6—fail to converge even after 500 relaxation steps (final Fmax 0.44 eV Å-1), indicating that CHGNet cannot locate a stable minimum for these chemistries. The As2In4Pd6 relaxation releases 15.4 eV/atom, roughly its DFT formation energy. These results confirm that CHGNet’s entire potential energy surface is fundamentally incorrect for these compositions, not merely its force predictions at a single probe geometry.
3.5 Failure analysis: UQ orthogonality and training-data gaps
A natural question is whether existing uncertainty quantification (UQ) methods can identify these blind spots without adversarial auditing. We test this by computing perturbation-based uncertainty—the approach used by CAGO11—for 200 WBM materials. For each material, we generate 5 structures with Gaussian perturbations ( Å) and measure CHGNet prediction variance.
The point-biserial correlation between perturbation uncertainty and compositional failure is (), indicating that structural sensitivity is orthogonal to compositional failure. Spearman rank correlation reveals a marginal monotonic relationship (, ), and failed materials show higher mean uncertainty (Cohen’s )—but these effects are too weak for reliable deployment screening. The 50 highest-uncertainty materials show a paradoxically negative correlation () with actual failure, meaning UQ-based screening would miss the most consequential blind spots. This directly challenges the assumption that model uncertainty20; 29; 30 can serve as a proxy for reliability: a model can be confidently wrong on entire chemical families. Adversarial compositional auditing and structural perturbation-based UQ capture independent failure dimensions and should be deployed jointly.
The blind spots are also not random with respect to the periodic table. Per-element analysis across 84 elements reveals that f-block elements have significantly higher DFT/CHGNet disagreement (Kruskal–Wallis ; Extended Data Fig. 5). Elements targeted by adversarial strategies show higher disagreement than non-targeted elements (59.9% vs 55.4%, Mann–Whitney ). The rarest frequency quintile contains 41% of adversarial elements; the most common contains none.
Cross-referencing 814 WBM materials against the JARVIS-DFT database39 (OptB88vdW functional) confirms 682 blind spots across DFT functionals, consistent with training data from the Materials Project42, OQMD40 and AFLOW41. Heavy elements show significantly higher disagreement (mean atomic mass , max ; Mann–Whitney).
3.6 Envelope refinement and formal certification
The refined envelope delineates reliable from unreliable MLIP predictions. As a concrete example: starting from an initial claim that CHGNet is reliable for , the heuristic adversary discovers Cu7Zn1 (brass) at step 47 as a counterexample with prediction error 10.49 eV atom-1. After 200 queries (86% counterexample rate, 61 unique materials), bootstrap refinement tightens the envelope to key bounds: (all counterexamples above this threshold failed), mean atomic mass 164 u, bandgap (PBE) 0.92 eV—consistent with known CHGNet limitations. Iterative refinement (4 rounds of attack–refine) compresses the safe envelope by 75–91% per feature dimension (Extended Data Fig. 4).
The envelope compiles into five Lean 4 proof modules ( lines, Lean 4.27.0): safety theorem with physical axioms, perturbation stability via interval arithmetic, error propagation composing DFT uncertainty (0.1 eV atom-1) with MLIP threshold via triangle inequality, monotone refinement soundness, and evaluation metric properties (CX rate monotonicity, CI convergence with budget). All axioms are explicit and inspectable. The proofs verify reasoning correctness—that conclusions follow from stated assumptions—analogous to AI-Descartes6 combining data and theory, and providing the machine-checkable evidence that structured safety cases16 currently lack.
3.7 Prospective validation
The preceding results establish that PCM discovers genuine blind spots. A stronger question is whether those discoveries generalise: can compositional features identified through adversarial auditing predict failures on materials the framework has never seen?
Protocol. We split the full 25,000-material WBM benchmark into a discovery set (15,000 materials, 60%) used for adversarial auditing and envelope refinement, and a held-out validation set (10,000 materials, 40%) never seen during any stage of the PCM pipeline. A gradient-boosted classifier (scikit-learn GradientBoostingClassifier, 5 random seeds) is trained to predict binary failure ( eV Å-1) from the 8 compositional features used throughout this work.
Results. The classifier achieves AUC-ROC on the held-out validation set (mean s.d. across 5 random seeds), demonstrating that PCM-discovered failure patterns are not overfit to the discovery set but capture genuine compositional regularities. Precision at 20% recall (P@20%) is 1.000: the top 20% of risk-ranked materials are 100% actual failures. This means a deployment pipeline that flags the highest-risk quintile achieves perfect precision with zero false alarms.
Key features. The top predictive features are n_sites (number of atoms per unit cell), volume_per_atom, and max_z (maximum atomic number)—all compositional descriptors that PCM counterexamples naturally concentrate on during adversarial search. These are simple, interpretable features that materials scientists can act on directly: compositions with many sites, large volumes per atom, or heavy elements warrant additional DFT validation before trusting MLIP predictions.
Implication. This result transforms PCM from a retrospective auditing tool into a predictive intervention: rather than auditing every material individually, a PCM-trained risk model can screen large candidate libraries and flag likely failures before expensive DFT computation. The combination of adversarial discovery (identifying where failures occur) and prospective prediction (generalising why they occur) closes the loop between falsification and deployment.
3.8 Cross-MLIP prospective transfer
A critical question for deployment is whether PCM-discovered failure patterns are specific to a single MLIP architecture or capture general vulnerability patterns shared across architectures. We test this directly: can a risk model trained on one MLIP’s failures predict failures of a different MLIP?
Protocol. Using a separate 10,000-material multi-model benchmark (CHGNet + MACE), we split into discovery (6,000, 60%) and validation (4,000, 40%) sets. A gradient-boosted classifier is trained on one model’s failures from the discovery set and evaluated on both models’ failures on the held-out validation set (5 random seeds).
Results. Same-model baselines confirm strong predictive power: CHGNetCHGNet AUC-ROC = , MACEMACE AUC-ROC = . Cross-architecture transfer retains substantial predictive signal: CHGNetMACE AUC-ROC = , MACECHGNet AUC-ROC = (average cross-MLIP AUC-ROC = ). Both cross-transfer directions exceed the 0.5 random baseline by a wide margin ( across all seeds), indicating that a significant fraction of MLIP vulnerability is compositionally determined rather than architecture-specific.
Feature importance overlap. The Pearson correlation between CHGNet and MACE feature importances is , with 2.2 out of 3 top features shared across models (consistently n_sites and volume). This high overlap confirms that both architectures struggle with the same compositional drivers—large unit cells and heavy elements—even though their absolute failure sets are largely disjoint ( for raw errors).
Consensus prediction. Training on either model and predicting “any model fails” yields AUC-ROC (CHGNet-trained) and (MACE-trained), suggesting that single-model PCM audits capture a substantial portion of the shared vulnerability landscape.
Implication. This result addresses the key generalizability concern: PCM features are not CHGNet-specific artifacts. A risk model trained on one MLIP’s failures provides actionable early warning for other architectures, with a shared compositional vulnerability signal () and architecture-specific residuals. This supports deployment of PCM as a partially architecture-transferable auditing framework, where cross-MLIP risk models provide useful triage and per-MLIP retraining yields optimal performance.
3.9 Pipeline impact
To quantify the practical consequences of MLIP blind spots, we evaluate single-model screening performance on the WBM benchmark.
Single-model screening. Using CHGNet as the sole stability filter on 25,000 WBM materials (the standard high-throughput screening protocol), precision (fraction of CHGNet-trusted materials that are truly DFT-stable) is 0.47 and recall (fraction of DFT-stable materials that CHGNet identifies) is 0.07. This means that (a) roughly half of CHGNet-predicted “stable” materials are actually DFT-unstable (false positives), and (b) CHGNet misses 93.0% of DFT-stable materials (false negatives; distinct from the 93.2% counterexample rate on 5K, which counts all failure types). The false negative rate is the more consequential metric: a screening pipeline that rejects 93% of genuinely stable materials is systematically excluding the materials it aims to find.
PCM-audited screening. Applying the PCM-refined envelope as a pre-filter—flagging materials in compositional regions where CHGNet is known to be unreliable—reduces the false positive rate from 10.7% to 8.1% (a 24% relative reduction). Materials flagged by the PCM envelope are routed to DFT validation rather than trusted at MLIP-level, ensuring that blind-spot regions receive appropriate scrutiny. The prospective risk model (Section 3.7) further enables prioritised DFT allocation: the highest-risk quintile can be validated first, maximising the yield of genuine discoveries per DFT calculation.
These numbers establish the pipeline impact of unaudited MLIP deployment: without adversarial auditing, a CHGNet-based screening pipeline operates at 7% recall for DFT-stable materials, rejecting 93.0% of the materials it should identify as promising candidates.
3.10 Multi-MLIP deployment simulation
The preceding sections establish that (a) blind spots are architecture-specific and (b) PCM risk models transfer across architectures. A practitioner’s natural question is: how should I change my screening pipeline tomorrow? We simulate five deployment strategies on a 10,000-material subset of WBM with both CHGNet and MACE predictions (distinct from the 5K single-model and 25K prospective benchmarks).
Screening strategies. Using force reliability ( eV Å-1) as the trust criterion:
CHGNet-only: precision 0.564, recall 0.676, false negative rate (FNR) 32.4% (6,829 trusted)
MACE-only: precision 0.529, recall 0.174, FNR 82.5% (1,882 trusted)
Union (trust if either reliable): precision 0.559, recall 0.698, FNR 30.2% (7,116 trusted)
Consensus (trust only if both reliable): precision 0.545, recall 0.152, FNR 84.8% (1,595 trusted)
Consensus screening is too conservative, missing 84.8% of DFT-stable materials. Union achieves the best no-DFT recall (69.8%), improving over CHGNet-only by recovering materials where MACE is reliable but CHGNet is not (and vice versa). Note that the Union strategy is effectively an ensemble disagreement baseline—trusting any material where at least one model is confident. Even this natural multi-model heuristic achieves only 69.8% recall; adding the PCM risk model raises recall to 78.0%, demonstrating that adversarial compositional auditing captures failure patterns beyond what ensemble agreement alone can detect.
PCM-audited screening. Adding the PCM risk model to CHGNet screening (routing the top 20% risk-ranked materials to DFT) raises screening precision (fraction of trusted materials that are DFT-stable) from and recall (fraction of DFT-stable materials correctly identified) from —recovering 10.4% more DFT-stable materials that unaudited CHGNet would reject. Materials in PCM-flagged regions are DFT-verified rather than blindly trusted or rejected.
DFT budget efficiency. The PCM risk model prioritises which materials most need DFT validation. At a 20% DFT budget (800 materials out of 4,000 in the validation set), PCM-ranked allocation achieves a DFT yield of (75.6% of DFT-evaluated materials are genuinely stable), compared to for random DFT allocation—a 34% improvement in DFT efficiency. This means each DFT calculation is 1.34 more likely to discover a stable material when guided by PCM risk ranking.
Blind spot coverage. 72.1% of all force failures (2,818 / 3,909) are model-specific: 1,409 fail CHGNet only, 1,409 fail MACE only. Among the 5,703 DFT-stable materials, 42.7% are flagged as unreliable by at least one MLIP—yet only 13.0% are flagged by both. This confirms that multi-MLIP auditing catches blind spots invisible to any single model.
Practitioner recommendation. Based on these results, the optimal deployment strategy is: (1) screen with union of available MLIPs to maximise recall, (2) apply PCM risk model to flag compositional regions of concern, and (3) allocate DFT budget preferentially to PCM-flagged materials. This three-step protocol improves screening recall by 10% absolute and DFT efficiency by 34% relative, at negligible computational cost beyond the initial PCM audit.
3.11 Deployment case study: thermoelectric screening
To demonstrate the practical impact of PCM-audited screening in a realistic discovery workflow, we simulate a thermoelectric materials campaign. A researcher screens 647 WBM candidates with band gaps in the thermoelectric-relevant range (0.1–1.5 eV), of which 488 are DFT-stable (ground truth unknown to the screener).
Head-to-head comparison. Five protocols are evaluated (Table 2):
| Protocol | Discov. | Missed | False leads | Precision | DFT evals |
| CHGNet-only | 244 | 244 (50.0%) | 63 | 0.795 | 0 |
| MACE-only | 63 | 425 (87.1%) | 14 | 0.818 | 0 |
| Union (no DFT) | 253 | 235 (48.2%) | 65 | 0.796 | 0 |
| Union + random DFT | 296 | 192 (39.3%) | 52 | 0.850 | 130 |
| PCM-audited (ours) | 306 | 182 (37.3%) | 53 | 0.852 | 130 |
Key result. The PCM-audited protocol discovers 62 additional stable thermoelectrics that CHGNet-only screening would miss entirely—a 25.4% improvement in discovery yield. The miss rate drops from 50.0% to 37.3%, while precision improves from 0.795 to 0.852 (fewer false leads). This requires 130 DFT evaluations (20.1% of candidates), targeted by the PCM risk model at the highest-risk compositions. Compared to random DFT allocation at the same budget, PCM-ranked allocation recovers 53 discoveries versus 43 for random—a 23% improvement in DFT targeting efficiency.
Recovered materials. Among the discoveries recovered by the PCM-audited protocol are WSe2 (gap = 1.47 eV, eV/atom), IrSbZr (gap = 1.48 eV, half-Heusler thermoelectric), and BaBiScO6 (gap = 1.46 eV, double perovskite)—all DFT-stable materials that CHGNet flags as unreliable (max force 50 eV Å-1) due to architecture-specific blind spots. These are precisely the materials a high-throughput pipeline aims to discover.
Practical implication. For a research group screening 650 thermoelectric candidates, the PCM-audited protocol adds 60 genuine discoveries at the cost of 130 DFT calculations—roughly 0.5 DFT per additional discovery. Given that a single DFT relaxation costs minutes to hours, this represents a favourable trade-off: each additional DFT calculation targeted by PCM is 1.23 more likely to yield a discovery than random allocation.
3.12 Cross-domain generalizability
To test whether the PCM framework generalises beyond solid-state materials, we apply the identical pipeline to three independent domains—molecular property prediction (QM9), drug solubility (ESOL), and tabular ML regression (California Housing)—running extended experiments across all three: multi-strategy comparison (5 strategies 3 seeds per domain), budget sensitivity (5 budgets 3 strategies 3 seeds per domain), full pipeline verification, and failure characterisation. These experiments test framework generalization, not merely breadth: each domain has different data modalities, model architectures, and failure patterns, yet the same three-stage pipeline (falsify, refine, certify) applies without modification.
Molecular property prediction (QM9). Two real graph neural networks—SchNet52 (455K parameters, distance-based message passing) and DimeNet++53 (1.9M parameters, directional message passing)—were trained for 200 epochs on 8,000 QM9 molecules (2,000 held-out test set, 10,000 total)49 on an A4000 GPU. SchNet achieves test MAE = 0.125 eV (median error 0.101 eV); DimeNet++ achieves test MAE = 0.087 eV (median error 0.064 eV). The models produce moderate error correlations (Pearson ) with architecture-specific failure modes: at a 1.0 eV threshold, 6 molecules fail DimeNet++ only and 0 fail SchNet only—confirming disjoint blind spots even for well-trained GNNs. Top failure predictors are n_atoms and molecular_weight. At budget = 200, grid sampling achieves the highest CX rate (), followed by sobol (), LHS (), and random (), while the heuristic adversary discovers 23/200 counterexamples (). The heuristic’s lower rate reflects an informative mismatch: the stress model targets large, polar molecules, but real SchNet/DimeNet++ errors are not concentrated at stress extremes—uniform samplers cover the input space more broadly and find more failures spread across the space.
Drug solubility prediction (ESOL). A Morgan fingerprint Random Forest and a descriptor-based neural network predict aqueous solubility on 1,128 molecules (Delaney 2004). Failures are largely disjoint: 334 molecules fail the fingerprint model only, 58 fail the neural network only, and just 86 fail both (error correlation ). The heuristic adversary achieves CX rate versus grid , LHS , random and sobol . LogP is the strongest predictor of fingerprint model failure, while heavy_atoms drives neural network failure—again, distinct failure predictors for each architecture.
Tabular ML regression (California Housing). A GradientBoosting tree and an MLP regressor predict median house values on 20,640 California census tracts, demonstrating PCM in a genuinely non-chemistry domain. The GBT achieves MAE 0.31 ($31K) and the MLP MAE 0.34 ($34K). Error correlation is high (), with top failure predictors ave_occup, med_inc, and population. At budget = 200, grid achieves the highest CX rate (), followed by sobol (), random (), LHS (), and heuristic (). The heuristic adversary discovers 24/200 counterexamples (12.0%), with meaningful envelope refinement and bootstrap CIs computed. Lean 4 proofs verify for the tabular domain, confirming formal certification generalises beyond chemistry entirely.
Budget sensitivity. CX rate saturates early in all three domains. For QM9, the heuristic adversary is stable at – across all budgets while random saturates at budget = 100 (36.3%) and sobol at budget = 200 (37.8%). For ESOL, heuristic saturates at budget = 200 (80.3%), random at budget = 500 (59.4%), and sobol remains stable at –. For Tabular, heuristic saturates at budget = 100 (15.3%), random at budget = 200 (40.0%), and sobol at . This confirms the WBM finding (Section 5) that a modest adversarial budget suffices for failure discovery across diverse ML domains.
Full pipeline. All three domains complete the entire PCM pipeline—attack, envelope refinement, Lean 4 proof generation—with machine-checked proofs verified by lake build in all cases (QM9: 23/200 CX, 11.5%; ESOL: 161/200 CX, 80.5%; Tabular: 24/200 CX, 12.0%). This provides strong evidence that formal certification of ML reliability claims extends beyond solid-state materials to molecular property prediction with real GNNs and non-chemistry tabular ML alike.
All five experiments demonstrate that architecture-specific blind spots, adversarial auditing, multi-strategy comparison, and formal verification generalise across chemistry domains with auditable ML models; the tabular ML experiment confirms PCM generalises to non-chemistry domains.
4 Discussion
Prospective validation as the strongest evidence. The most significant result is not the discovery of blind spots—which prior work has noted qualitatively—but the demonstration that PCM-discovered features predict failures on unseen materials (AUC-ROC ; P@20% = 1.000, i.e. perfect precision among the top-risk quintile with zero false alarms). This transforms adversarial auditing from a retrospective exercise into a prospective tool. The key insight is that compositional features concentrating counterexamples (n_sites, volume_per_atom, max_z) generalise to unseen chemistries—a property that conformal prediction (marginal coverage only) and perturbation-based UQ () do not provide.
Deployment impact. Single-model screening misses 93% of DFT-stable materials; multi-MLIP screening with PCM risk-ranked DFT allocation improves efficiency by 34%. The thermoelectric case study (Section 3.11) demonstrates that these improvements translate to concrete discovery gains: 62 additional stable materials recovered at a cost of 130 targeted DFT calculations. The recommended three-step protocol—union screening, PCM risk flagging, prioritised DFT—is immediately adoptable. Critically, 72.1% of failures are model-specific, confirming that no single-model audit can substitute for multi-MLIP coverage.
Formal verification and the specification gap. The Lean 4 proofs are explicitly conditional: every approximation (DFT at 0.1 eV atom-1, MLIP threshold at 2.0 eV atom-1) appears as an inspectable axiom with constructive non-vacuity witnesses (Extended Data Fig. Extended Data Fig. 9). This addresses the specification gap17 by making assumptions explicit rather than hiding them in implementation details. Combined with prospective failure prediction, this constitutes a new approach to MLIP validation: traditional benchmarks provide aggregate scores; conformal prediction provides coverage guarantees without compositional specificity; PCM provides falsifiable safety certificates with both formal guarantees and predictive power. Cross-MLIP transfer (AUC-ROC cross-model, feature importance ) confirms that vulnerability patterns are partially shared across architectures.
Cross-domain generalizability. Extended experiments across three chemistry domains (WBM, QM9 with real SchNet/DimeNet++ GNNs, ESOL) and one non-chemistry domain (California Housing, tabular ML) confirm that the three-stage pipeline generalises across data modalities and model architectures (Section 3.12). Each domain exhibits distinct failure predictors per architecture, and all produce verified Lean 4 proofs.
Accessibility and practical considerations. The full multi-strategy, multi-MLIP audit costs $18.13; a single-model audit with purely algorithmic adversaries costs nothing. CX rate saturates by budget = 50, though discovery diversity grows linearly with budget (24 unique materials at budget = 25 vs 231 at budget = 500). PCM selects adversarial points for auditing (200 queries), not retraining (100K structures)—at budget = 200, the probability of missing a failure mode of prevalence 5% is 0.003, providing a PAC-style coverage guarantee without model modification. A PCM audit report provides four actionable outputs: (1) a pass/fail decision with quantified confidence, (2) an explicit compositional map of unreliable regions, (3) machine-checkable proofs, and (4) a prospective risk model for screening unseen materials.
| Step | Action |
| 1. Define | Write a YAML envelope specifying compositional bounds and failure thresholds. |
| Example: , prediction error 2 eV atom-1, force 50 eV Å-1. | |
| 2. Attack | Run adversarial search: python -m pcm attack --model <yours> --budget 200. |
| PCM deploys 6 strategies (random, heuristic, grid, LHS, Sobol, LLM) to find counterexamples. | |
| 3. Certify | Inspect the refined envelope and machine-checked Lean 4 proofs. |
| The proofs certify: if the assumptions hold, then the safety claim follows. |
Limitations. (1) Ground truth is computational DFT, not experiment—though independent QE recomputation (20/20 converging on adversarially selected compositions), cross-functional (JARVIS) validation, and literature-verified cases (TlBiSe2) provide three independent lines of evidence. (2) The WBM oracle is database retrieval, not structure generation; live CHGNet evaluation on 50 approximate (golden-ratio) structures shows moderate energy correlation (, ) and force correlation (, ) with database values. This gap arises because the database uses DFT-relaxed equilibrium structures while our approximate structures use simple cubic cells reconstructed from formula and volume. The blind-spot classification itself, however, derives from the WBM benchmark where both DFT and CHGNet evaluate relaxed structures. (3) The benchmark’s high base CX rate means adversary advantage is in diversity, not rate. (4) Lean proofs verify reasoning correctness, not physics. (5) Cross-MLIP transfer (Section 3.8) shows moderate but imperfect generalisation (average AUC-ROC = ): while a CHGNet-trained risk model provides useful early warning for MACE failures, architecture-specific retraining improves AUC-ROC by , suggesting that optimal deployment uses per-MLIP auditing.
5 Methods
5.1 Framework and oracle models
PCM operates on three abstractions: Environment (compositional features: , , electronegativity, volume per atom, atomic mass), Design (structural parameters: , bandgap), and Outcomes (prediction error, maximum force). A safety claim asserts outcomes satisfy thresholds for all (env, design) within specified bounds. Six adversary strategies (random, heuristic, LHS, Sobol, grid, LLM) are compared at equal budget.
The WBM oracle maps adversary-proposed feature vectors to nearest real materials via KDTree lookup () and returns pre-computed CHGNet/DFT comparisons. TensorNet and MACE oracles use the same WBM feature space with live MLIP inference via matgl and mace-torch, respectively. Stability classification uses eV atom-1. KDTree-indexed feature lookup enables sub-millisecond queries at full WBM scale (0.43 ms per query at 257K vs 0.14 ms at 5K; coverage at budget = 1,000: 504 unique materials at 257K vs 347 at 5K), so the framework scales to production datasets without re-running MLIPs at attack time.
5.2 Adversary strategies
Six strategies are evaluated:
Random: Uniform sampling within envelope bounds.
Heuristic: Beta-distribution biased toward stress directions specified in envelope YAML.
Grid: Uniform grid over envelope bounds.
LHS: Latin Hypercube Sampling (scipy).
Sobol: Sobol quasi-random sequences (scipy).
LLM: Language model adversary receiving evaluation history and proposing numerical feature vectors targeting likely failure regions. On API failure, degrades to random sampling with exponential backoff. Temperature = 0.2 for all models.
5.3 Envelope refinement
Environment bounds are tightened to the 75th percentile of passing evaluations using the conservative 95% bootstrap CI (1,000 resamples). Minimum sample gate: 30 passing evaluations. Design requirements are stress-direction-aware.
5.4 Lean 4 certification
Refined envelopes compile into five proof modules: core safety theorem, perturbation stability, error propagation (composing DFT + MLIP uncertainty via triangle inequality), monotone refinement soundness, and evaluation metric properties (CX count monotonicity, CI width convergence, PAC coverage guarantee). Lean 4.27.05, no mathlib dependency. This builds on recent applications of theorem provers to scientific reasoning9; 18.
5.5 Prospective validation protocol
The WBM benchmark (25,000 materials) is split 60/40 into discovery (15,000) and validation (10,000) sets. The discovery set is used for all adversarial auditing and envelope refinement. A scikit-learn GradientBoostingClassifier (200 estimators, max depth 4, learning rate 0.05) is trained on the 8 compositional features (n_elements, max_z, mean_electronegativity, std_electronegativity, volume_per_atom, mean_atomic_mass, n_sites, bandgap_pbe) to predict binary failure ( eV Å-1). Results are aggregated across 5 independent random seeds (each with different discovery/validation splits). Performance is evaluated on the held-out validation set using AUC-ROC, precision-recall, and precision at % recall (P@%).
5.6 Benchmark construction
25,000 materials sampled from 257K WBM structures2, characterised by 8 compositional features with pre-computed CHGNet v0.3.0 predictions and DFT reference values. Full 257K available for scaling validation. Anomaly score:
robust to alternative weightings (Kendall across 5 configurations).
5.7 Independent DFT validation
Quantum ESPRESSO 6.7, projector augmented wave (PAW) Perdew–Burke–Ernzerhof (PBE) with Standard Solid-State Pseudopotentials (SSSP) v1.3.0 efficiency library, 60/480 Ry cutoffs, -mesh, cold smearing ( Ry), mixing , convergence threshold Ry. TlBiSe2 converges in 13 SCF steps (207.6 s). For the 20-material validation (Section 3.4), materials are selected from 811 candidates meeting: DFT-stable, CHGNet-unstable, max , . Structures use golden-ratio fractional coordinate spacing (; deterministic, non-overlapping) with DFT-derived cell volumes from WBM. Each SCF calculation runs on 8 CPU cores with a 3,600 s timeout. All 20 converge: 18 on first pass (47 min total) and 2 after retry with expanded cells and tighter mixing (additional 2 hours). Total first-pass wall time for all 20 attempts is 54 min including the 2 that initially failed.
5.8 Cross-MLIP evaluation
Three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1-PES via matgl 2.0.6, evaluated in float32 precision), and MACE-MP medium4 (via mace-torch, float64)—are evaluated on 5,000 WBM materials (random sample, seed = 42). Structures are built from WBM composition, volume, and Wyckoff symmetry data using golden-ratio fractional coordinate spacing with correct DFT-derived cell volumes. Six pairwise correlations (energy and force) are computed with bootstrap 95% CIs (1,000 resamples). Force failure threshold: 50 eV Å-1.
Data Availability
The WBM benchmark (25,000 materials), all experimental runs and generated Lean 4 proofs are included in the repository. The full WBM dataset (257K) is available from Matbench Discovery2.
Code Availability
PCM is open-source at https://github.com/abhinaba/alloy_pcm.
Reproduction: pip install -e . && python -m pcm demo.
Acknowledgements
We thank the developers of CHGNet, MACE and TensorNet for making their models publicly available, and the Matbench Discovery team for the WBM benchmark infrastructure.
Author Contributions
A.B. conceived the project, developed the PCM framework, performed all experiments and wrote the manuscript. P.C. supervised the project and reviewed the manuscript.
Competing Interests
The authors declare no competing interests.

### v2

@@
 1 Introduction
-Universal MLIPs—CHGNet1, MACE4, TensorNet27, ALIGNN35, SevenNet36, EquiformerV237—underpin high-throughput materials screening38, yet they are deployed without formal reliability guarantees.
-Matbench Discovery2 and earlier benchmarks10 evaluate 45+ models on 257K WBM (Wang–Botti–Marques3) structures, but aggregate accuracy metrics cannot answer the deployment-critical question: on which chemistries is this MLIP unreliable?
+Universal MLIPs—CHGNet1, MACE4, TensorNet27, ALIGNN35, SevenNet36, EquiformerV237—underpin high-throughput materials screening38; 55, yet they are deployed without formal reliability guarantees59; 60.
+Matbench Discovery2, MLIP Arena57, CHIPS-FF58 and earlier benchmarks10 evaluate 45+ models on 257K WBM (Wang–Botti–Marques3) structures, but aggregate accuracy metrics cannot answer the deployment-critical question: on which chemistries is this MLIP unreliable?
 This creates a quantifiable safety gap.
@@
 Conformal prediction19 provides distribution-free marginal coverage guarantees applied to interatomic potentials20.
-Monte Carlo dropout29, deep ensembles30 and chemical-space uncertainty estimation31; 32 have been applied to molecular property prediction.
+Monte Carlo dropout29, deep ensembles30 and chemical-space uncertainty estimation31; 32, committee disagreement63 and misspecification-aware regression61; 62 have been applied to molecular property prediction.
 Critically, these methods provide marginal coverage (“95% of all materials fall within the prediction interval”) but cannot answer the deployment question which specific compositions are unreliable and why.
@@
 Adversarial testing in ML33; 34 targets input perturbations; falsification of cyber-physical systems7; 8 provides compositional testing frameworks.
+Komendantskaya64 extends proof-carrying code to neuro-symbolic programs, and Kamran et al.65 apply it to LLM-generated code completions.
 PCM instantiates the Guaranteed Safe AI framework13 for materials: the MLIP is the world model, the envelope is the specification, and Lean 4 proofs are the verifier output—providing machine-checkable evidence that structured safety cases currently lack.
@@
 3.1 Architecture-specific failure profiles
-We evaluate three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1)27; 28, and MACE-MP medium4—on 5,000 WBM materials.
+We evaluate three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1)27; 28, and MACE-MP medium4; 54—on 5,000 WBM materials.
 Structures are synthetic compositional probes—atoms placed at irrational (golden-ratio) fractional coordinates to avoid symmetry artifacts, with DFT-derived cell volumes (Methods)—designed to test whether each MLIP has learned the correct energy landscape for a given chemistry.
@@
 Cross-MLIP transfer (AUC-ROC cross-model, feature importance ) confirms that vulnerability patterns are partially shared across architectures.
+The recent extension of proof-carrying code to neuro-symbolic programs64 reinforces this direction: as ML models become embedded in scientific workflows, machine-checkable certificates become essential infrastructure, not optional overhead.
+Emerging benchmarks and foundation models.
+Recent large-scale efforts—OMat2455 (118M DFT calculations), UMA56 (500M+ structures), and MatPES28—have dramatically expanded MLIP training data, while MLIP Arena57 and CHIPS-FF58 benchmark physics-aware properties beyond formation energy.
+These developments are complementary to PCM: larger training sets reduce but do not eliminate architecture-specific blind spots (as our cross-MLIP analysis with three models trained on overlapping Materials Project data demonstrates), and emerging benchmarks still evaluate aggregate accuracy rather than compositional specificity.
+PCM provides the missing layer—identifying which compositions remain unreliable after training on massive datasets, enabling targeted DFT validation where it matters most.
+Committee disagreement methods63 and misspecification-aware UQ61; 62 focus on per-prediction uncertainty but do not produce the explicit compositional failure maps or formal certificates that PCM provides.
+Robust training strategies66 and improved data generation67 address reliability from the training side; PCM addresses it from the deployment side—both are needed for trustworthy MLIP deployment.
+LLM-assisted materials design70 could further integrate with PCM by using adversarially discovered failure regions to guide generative exploration away from unreliable compositional space.
 Cross-domain generalizability.
@@
 Lean 4.27.05, no mathlib dependency.
-This builds on recent applications of theorem provers to scientific reasoning9; 18.
+This builds on recent applications of theorem provers to scientific reasoning9; 18; LLM-assisted proving68; 69 could further automate proof generation in future work.
 5.5 Prospective validation protocol
@@
 5.8 Cross-MLIP evaluation
-Three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1-PES via matgl 2.0.6, evaluated in float32 precision), and MACE-MP medium4 (via mace-torch, float64)—are evaluated on 5,000 WBM materials (random sample, seed = 42).
+Three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1-PES via matgl 2.0.6, evaluated in float32 precision), and MACE-MP medium4; 54 (via mace-torch, float64)—are evaluated on 5,000 WBM materials (random sample, seed = 42).
 Structures are built from WBM composition, volume, and Wyckoff symmetry data using golden-ratio fractional coordinate spacing with correct DFT-derived cell volumes.